OpenStudy (anonymous):

A die is rolled 8 times. Find the probability. P(getting at least 7 even numbers)

OpenStudy (anonymous):

you roll 8 dice, the probability of you rolling an even number is 1/2 P(at least 7 evens) = P(7 even numbers) + P(8 even numbers) P(7 even numbers) = (1/2)^7 * (1/2) P(8 even numbers) = (1/2)^8 P(7 even numbers) + P(8 even numbers) = (1/2)^7 * (1/2) + (1/2)^8 (1/2)^8 + (1/2)^8 2*(1/2)^8 (1/2)^7